OpenStudy (anonymous):

Which sequences are geometric? Check all that apply. –2, –4, –6, –8, –10, … 16, –8, 4, –2, 1 –15, –18, –21.6, –25.92, –31.104, … 4, 10.5, 17, 23.5, 30, … 625, 125, 25, 5, 1, …

OpenStudy (defiance):

@apalmer004 The last question you asked on the others, Number 2 was not one of them, only the last 2. Sorry.

OpenStudy (misty1212):

weird i swear i just saw this exact question

OpenStudy (defiance):

Misty will be able to help you here, I gotta go to the movies

OpenStudy (misty1212):

what are you going to see?

OpenStudy (anonymous):

do you know the answer? @misty1212

OpenStudy (misty1212):

see which ones have a "common ratio" i.e if you divide one term by the previous one, you always get the same number

OpenStudy (misty1212):

not \[–2, –4, –6, –8, –10, …\] since there you are subtracting 2 each time, so it is "arithmetic"

OpenStudy (misty1212):

\[16, –8, 4, –2, 1\] is geometric since \[\frac{-8}{16}=\frac{8}{-8}=\frac{-2}{4}=\frac{-1}{-2}\] all the ratios are \(-\frac{1}{2}\)

OpenStudy (misty1212):

typo there, i meant \[\frac{-8}{16}=\frac{4}{-8}=\frac{-2}{4}=\frac{-1}{-2}\]

OpenStudy (misty1212):

for \[–15, –18, –21.6, –25.92, –31.104, … \] check using a calculator if \[\frac{-18}{-15}=\frac{-21.6}{-18}=\frac{-25.92}{-21.6}\]